# SFFT-based Homogenization: Using Tensor Trains to Enhance FFT-Based Homogenization

### v1

@@ Appendix F Discretization Scheme
Before investigating the various geometries, it is essential to determine the most appropriate discretization scheme by analyzing the Green-Eshelby Operator. This decision is not straightforward, as the iteration time depends on multiple factors: the smoothness of the solution field (and consequently its rank), as well as the rank of the Green-Eshelby Operator itself. While more complex discretization schemes may result in smoother solution fields, they could also lead to a higher rank in the Green-Eshelby Operator, potentially impacting the efficiency of the algorithm.
@@
Interestingly, Figure 17 b) reveals that the Green-Eshelby Operator in the centered grid scheme actually has a lower rank than the staggered grid over the range of discretization orders $n$ considered. However, the staggered grid scheme seems to eventually surpasses the centered grid at higher discretization orders due to its improved scaling behaviour.

### v3

@@ Appendix G Discretization Scheme
Before investigating the various geometries, it is essential to determine the most appropriate discretization scheme by analysing the Green-Eshelby Operator. This decision is not straightforward, as the iteration time depends on multiple factors: the smoothness of the solution field (and consequently its rank), as well as the rank of the Green-Eshelby Operator itself. While more complex discretization schemes may result in smoother solution fields, they could also lead to a higher rank in the Green-Eshelby Operator, potentially impacting the efficiency of the algorithm.
@@
Figure 17 b) shows that the Green-Eshelby Operator in the centered grid scheme has a lower rank than in the staggered grid for the considered discretization range. However, at higher discretization orders, the staggered grid surpasses the centered grid due to its improved scaling behaviour.
The iteration times and strain rank growth in the centered scheme reveal an interesting phenomenon: hourglassing. This numerical artifact introduces spurious, low-frequency oscillations that do not correspond to real physical deformations, increasing the strain rank. As a result, iteration times rise despite the centered grid having the lowest Green-Eshelby operator rank.